# iA: Imperative Learning-based A Search for Path Planning

### v2

@@ Abstract
Path planning, which aims to find a collision-free path between two locations, is critical for numerous applications ranging from mobile robots to self-driving vehicles. Traditional search-based methods like A search guarantee path optimality but are often computationally expensive when handling large-scale maps. While learning-based methods alleviate this issue by incorporating learned constraints into their search procedures, they often face challenges like overfitting and reliance on extensive labeled datasets. To address these limitations, we propose Imperative A (iA), a novel self-supervised path planning framework leveraging bilevel optimization (BLO) and imperative learning (IL). The iA framework integrates a neural network that predicts node costs with a differentiable A search mechanism, enabling efficient self-supervised training via bilevel optimization. This integration significantly enhances the balance between search efficiency and path optimality while improving generalization to previously unseen maps. Extensive experiments demonstrate that iA outperforms both classical and supervised learning-based methods, achieving an average reduction of 9.6% in search area and 15.2% in runtime, underscoring its effectiveness in robot path planning tasks.
@@ I INTRODUCTION
Path planning, determining a route from a starting position to a destination in a given environment, is one of the most basic research topics in robot navigation[1]. It is critical for various applications such as those requiring the shortest distance, minimal energy use, or maximum safety [2].
@@
Due to their optimality, search-based algorithms, such as Dijkstra’s algorithm [3] and A [4], have become one of the most popular methods in path planning [5]. However, a major drawback of search-based methods is their inefficiency for large-scale maps, since their search space and runtime complexity grow exponentially with the increasing size of maps [6]. Although heuristic strategies such as D [7] have been introduced to guide more efficient searching directions, their onboard real-time performance is still unsatisfactory, especially on low-power devices and mobile robots.
@@
Figure 1: The comparison of the searched area of search-based path planning methods. (a), (b), and (c) display the planning results by using A, Neural A, and iA (ours), where the red and green circles represent the start node and goal node, while the red pixels and green pixels represent the generated path and search area.
@@
Recently, researchers have adopted data-driven and supervised learning approaches, such as Neural A [8] and TransPath [9], to predict constrained search space. While these methods lose theoretical optimality guarantees, they have garnered increasing attention due to their ability to significantly reduce the search area and efficiently find near-optimal paths with only a slight increase in path length. However, as they rely on supervised learning, they require extensive pre-processing to get the optimal paths in each training iteration. Moreover, they are prone to overfitting, limiting their effectiveness in out-of-training maps [5].
@@
To eliminate the requirement for pre-processing and overcome the overfitting challenge, we introduce bilevel optimization (BLO) into search-based path planning, inspired by imperative learning (IL) [10]. This results in a novel self-supervised path planning framework, which we refer to as imperative A (iA). Specifically, iA consists of an upper-level optimization, which is a neural network to predict a cost for each node, and a lower-level optimization, which is a differentiable A search engine. Different from supervised learning methods like Neural A [8], the BLO in iA enables self-supervised learning by co-optimizing the neural network and the A search. More importantly, iA enables a more accurate training direction for the neural network by computing a more accurate gradient back-propagating through the bilevel models. This leads to a significantly better generalization performance, as demonstrated in both theoretical analysis and practical validation. In summary, our contributions are shown as follows:
@@
  • We introduce a self-supervised data-driven framework, iA, for global path planning through bilevel optimization based on imperative learning. iA enables efficient near-optimal solutions while balancing search area and path length without overfitting via a differentiable self-supervision in the low-level optimization.
@@
  • We benchmark the performance of various classical and data-driven path planning methods against our proposed iA framework through extensive experiments on public datasets and simulation environments. The experimental results show that iA achieves an average of 11.0% better trade-off between path length and search area, while saving 15.2% runtime and 9.6% search areas than supervised learning methods for unseen maps, demonstrating its significantly better generalization abilities.
@@ II-A Classical Path Planning
The path planning problem, finding a collision-free path from the start node to the goal node with a given map, has developed for many years. In the beginning, the classical works solve the path planning problem by graph-search approaches [11], which iteratively search the nodes from the start node to the goal node, such as Breadth First Search (BFS). After that, some algorithms, such as Greedy BFS [12], A [4], and D [13], introduce the heuristic information and provide a search direction to improve the planning efficiency. These heuristic search-based algorithms can find the optimal path when their suitable heuristic functions are admissible. However, these search-based methods face the problem of low search efficiency with the increasing size of maps due to their exponential time complexity growth.
Unlike search-based pathfinding methods, sampling-based methods sample nodes and connect nodes to find a solution in the given map. As the number of sampling points increases, the solution is closer to the optimal path. Typically, they stop their sampling when finding a near-optimal path acceptable in the given task to improve their search efficiency. The sampling-based methods can alleviate the computational growth by increasing the dimension of maps. By using a tree structure to store nodes and establish connections between the sampled points, some methods have outstanding performance, such as RRT [14] and its derivatives [15, 16, 17, 18]. Besides, some sampling-based path planning methods [19, 20, 21] utilize the probabilistic road map to get solutions. However, these sampling-based methods can only get the near-optimal solution and may fail to find a path in complex environments with high obstacle density.
@@ II-B Learning-based Path Planning
Recently, with the development of deep learning technology, plenty of learning-based path planning methods have appeared. Some of them utilize the networks to extract environmental information and provide guidance for improving search efficiency. SAIL [22] aims to find a suitable heuristic function that could observe and process the whole map instead of the local map during planning. Besides, [23] and TransPath [9] integrate the U-net-based model [24] and Transformer-based model [25] to learn their heuristic functions, respectively. Some methods [26, 8] utilize their well-trained network to encode the original map to generate a guidance map. Then, the guidance map allows them to reduce their search area and improve their search efficiency. MPNet[27] uses two networks, Encoder net (Enet) and Planning net (Pnet), for the path planning task, where Enet handles the environment information, and Pnet can provide the next state of the robot under the current situation.
Some researchers use reinforcement learning (RL) [28] to solve the path planning task. They aim to find an end-to-end solution for the path planning task without any policies. They formulate the path planning problem as a Markov decision process where the next state in the path depends on the current state and surroundings. [29] demonstrates that it is possible to solve the path planning problem in grid maps using the RL-based method. [30] combines the path graph with Q-learning, efficiently finding the optimal path. However, these learning-based methods face problems of lacking interpretability and extensive data labeling.
@@ II-C Imperative Learning
Imperative learning (IL) [31] is an emerging self-supervised neuro-symbolic framework based on bilevel optimization, where the upper-level optimization is a data-driven model and the lower-level optimization is an interpretable reasoning engine. IL adopts a self-supervised manner. The results from lower-level optimization can provide an adaptive and precise gradient for the training of the data-driven model within upper-level optimization. For example, iPlanner [32] utilizes a B-spline interpolation algorithm to guide the networking training, generating safe and smooth trajectories with depth images. iSLAM [10] adopts the IL framework to integrate the front-end and back-end of visual odometry. The pose estimation guides the training to generate the poses aligned with geometrical reality. iMatching [33] utilizes the bundle adjustment module to train a neural model, generating the optimal translation matrix for pose estimation. iKap [34] leverages a differentiable model predictive control to achieve a kinematics-aware vision-to-planning framework. These works demonstrate the efficiency of IL in robot navigation tasks. To the best of our knowledge, iA is the first method using the IL to solve global path planning problems.
@@ III Methodology
Traditional search-based methods can guarantee finding the optimal paths, while deep neural networks have the ability to reduce the search area by extracting the map information and making predictions. To leverage the advantages of both learning-based and search-based path planning methods, the proposed imperative A (iA) adopts the imperative-learning (IL) strategy, including learning-based upper-level optimization and A based lower-level optimization.
@@ III-A Framework
The framework of iA is depicted in Fig. 2, which takes a planning instance as input, a three-layer tensor containing the information of obstacles, start, and goal positions, and outputs the near-optimal path (red) and search area (green). The iA framework can be formulated as a bilevel optimization:
@@
Figure 2: The framework of iA. The input is a three-layer tensor containing map information, start node, and goal node, and the output is the path with the search area. It mainly has three parts: the Instance Encoder, the Memory Storage, and the A search.
The main difference between previous methods, like Neural A [8], and iA is that the bilevel optimization (1) incorporates a metric-based optimization to provide a more accurate training direction for the network $f$, leading to a more “generalizable” smaller search area. Intuitively, if the lower-level optimization (1b) is removed and $\mu^{*}$ represents constant optimal paths (labels), then bilevel optimization (1) degrades to a single-level optimization and iA becomes Neural A, and then the network gradients $\nabla_{\theta}\mathcal{U}$ can be calculated as
@@
$\nabla_{\theta}U=\frac{\partial U}{\partial f_{\theta}}\frac{\partial f_{\theta}}{\partial\theta}+\boxed{\frac{\partial U}{\partial\mu^{*}}\frac{\partial\mu^{*}}{\partial L}(\frac{\partial L\partial f}{\partial f\partial\theta}+\frac{\partial L}{\partial\mu}+\frac{\partial L}{\partial M})},$ (3)
[…]
@@ III-B Lower-level Optimization
The lower-level optimization module is designed to achieve two objectives: (1) executing the searching process for path planning, and (2) providing differentiable computational gradients for back-propagation in upper-level network training. The classical A algorithm inherently relies on a discrete graph structure, presenting challenges in effectively providing differentiable losses for upper-level optimization. To overcome this critical challenge, the lower-level optimization employs a differentiable A (dA[8] as its search engine, back-propagating the loss of the path planning process.
In dA, all variables and parameters are represented in matrix formulations. Specifically, denote the open-list matrix $O$ and closed-list matrix $C$, which are binary and of the same size as the given map, to record the candidate nodes under evaluation and the nodes that have already been evaluated, respectively. The start node $n_{s}$, the goal node $n_{g}$, and the currently selected node $n^{\prime}$ are separately represented by one-hot indicator matrices $N_{s}$, $N_{g}$, and $N^{\prime}$, where only a single element is set to $1$, and all others are 00. Denote $s(n)$ as the accumulated path length from the start node and $h(n)$ as the estimated heuristic distance to the goal. The matrices $S$ and $H$ store the $s$ and $h$ values of each node. During the path planning process, dA iteratively selects node $n^{\prime}$ with the minimum $s(n)+h(n)$ from the open list $\mathcal{O}$ and updates the $S$ matrix accordingly, until the optimal path is found.
In Neural A, the neural network estimates the initial $s$ values for each node to accelerate the searching process. However, this manner results in error propagation: inaccuracies in $s(n)$ estimation during earlier iterations accumulate and affect subsequent search as the $s$ values are updated relying on the prior $s$. To eliminate this problem, we introduce a predication factor $p$ in iA to independently represent the influence of upper-level optimization in the Node Selection part of dA. $P$ is the matrix storing $p$ values of each node. Then, we iteratively select the node with minimum $s(n)+h(n)+p(n)$ in the $\mathcal{O}$. It is worth noting that if the $p(n)$ equals $(\omega-1)h(n)$, our A search algorithm equals weighted A search [35], where the $\omega$ is a pre-set weight factor amplifying the influence of the heuristic function. In iA, the node selection is formulated as:
@@
where $A\odot B$ is Hadamard product [36], $\tau$ is an empirically defined parameter, $\mathcal{L}_{max}(A)$ is the function to obtain the $argmax$ index matrix of $A$, and $P$ is the predicted cost by the instances encoder $f$ within upper-level optimization.
@@ III-C Upper-level Optimization
The upper-level optimization aims to increase the search efficiency and keep the sub-optimality. This objective also balances the search area and path length. We take advantage of neural network in the upper-level optimization because of its great power in information extraction and prediction. The neural model is a widely-used U-Net [24], where the decoder is used for feature extraction and the encoder for prediction. The usage of U-Net ensures that the shape of the predicted $P$ is consistent with the shape of the input maps, avoiding changing the structure of the network and retraining in new environments with different shapes, and providing convenience in robot applications.
Figure 3: The calculation of the path length.
@@
where $U(\mu)$ is the target function of upper-level optimization given $\mu$, $\mathcal{L}_{a}(\mu)$, $\mathcal{L}_{l}(\mu)$ are the search area loss and path length loss of path $\mu$, and $w_{a}$ and $w_{l}$ are the weights to adjust the ratio of the losses. We define $\mathcal{L}_{a}(\mu)$ as the number of visited nodes in one single path planning process, which is equivalent to counting the $1$ value in $C_{\mu}$:
$\mathcal{L}_{a}(\mu)=\sum C_{\mu},$ (6)
@@
In iA, the upper-level optimization updates parameters in the embedded network to predict the $P$, pursuing the balance between search area and path length while maintaining near-optimal path quality. Unlike the other data-driven methods [8, 9], directly calculating loss with constant optimal paths, our method dynamically calculates loss from the path planning results of lower-level optimization in each iteration. Thus, the network parameters are updated to find the optimal balance between search area and path length under temporary conditions, instead of invariably following the constant optimal paths. To validate this bilevel training process, we compare the effects of the conventional optimal-path-based training process with ours in the ablation study.

### v4

# iA*: Imperative Learning-based A* Search for Path Planning

@@ Abstract
Path planning, which aims to find a collision-free path between two locations, is critical for numerous applications ranging from mobile robots to self-driving vehicles. Traditional search-based methods like A* search guarantee path optimality but are often computationally expensive when handling large-scale maps. While learning-based methods alleviate this issue by incorporating learned constraints into their search procedures, they often face challenges like overfitting and reliance on extensive labeled datasets. To address these limitations, we propose Imperative A* (iA*), a novel self-supervised path planning framework leveraging bilevel optimization (BLO) and imperative learning (IL). The iA* framework integrates a neural network that predicts node costs with a differentiable A* search mechanism, enabling efficient self-supervised training via bilevel optimization. This integration significantly enhances the balance between search efficiency and path optimality while improving generalization to previously unseen maps. Extensive experiments demonstrate that iA* outperforms both classical and supervised learning-based methods, achieving an average reduction of 65.7% in search area and 54.4% in runtime, underscoring its effectiveness in robot path planning tasks.
@@ I INTRODUCTION
Path planning, determining a route from a starting position to a destination in a given environment, is one of the most basic research topics in robot navigation[review]. It is critical for various applications such as those requiring the shortest distance, minimal energy use, or maximum safety [Quan2020Survey].
@@
Due to their optimality, search-based algorithms, such as Dijkstra’s algorithm [dijkstra1959note] and A* [hart1968formal], have become one of the most popular methods in path planning [gonzalez2015review]. However, a major drawback of search-based methods is their inefficiency for large-scale maps, since their search space and runtime complexity grow exponentially with the increasing size of maps [russell2016artificial]. Although heuristic strategies such as D* [dstar] have been introduced to guide more efficient searching directions, their onboard real-time performance is still unsatisfactory, especially on low-power devices and mobile robots.
@@
Figure 1: The comparison of the searched area of search-based path planning methods. (a), (b), and (c) display the planning results by using A*, Neural A*, and iA* (ours), where the red and green circles represent the start node and goal node, while the red pixels and green pixels represent the generated path and search area.
@@
Recently, researchers have adopted data-driven and supervised learning approaches, such as Neural A* [yonetani2021nastar] and TransPath [kirilenko2023transpath], to predict constrained search space. While these methods lose theoretical optimality guarantees, they have garnered increasing attention due to their ability to significantly reduce the search area and efficiently find near-optimal paths with only a slight increase in path length. However, as they rely on supervised learning, they require extensive pre-processing to get the optimal paths in each training iteration. Moreover, they are prone to overfitting, limiting their effectiveness in out-of-training maps [gonzalez2015review].
@@
To eliminate the requirement for pre-processing and overcome the overfitting challenge, we introduce bilevel optimization (BLO) into search-based path planning, inspired by imperative learning (IL) [fu2023islam]. This results in a novel self-supervised path planning framework, which we refer to as imperative A* (iA*). Specifically, iA* consists of an upper-level optimization, which is a neural network to predict a cost for each node, and a lower-level optimization, which is a differentiable A* search engine. Different from supervised learning methods like Neural A* [yonetani2021nastar], the BLO in iA* enables self-supervised learning by co-optimizing the neural network and the A* search.  Fig. 1 presents a visual comparison between Neural A* and iA*. More importantly, iA* enables a more accurate training direction for the neural network by computing a more accurate gradient back-propagating through the bilevel models. This leads to a significantly better generalization performance, as demonstrated in both theoretical analysis and practical validation. In summary, our contributions are shown as follows:
@@
  • We introduce a self-supervised data-driven framework, iA*, for global path planning through bilevel optimization based on imperative learning. iA* enables efficient near-optimal solutions while balancing search area and path length without overfitting via a differentiable self-supervision in the low-level optimization.
@@
  • We benchmark the performance of various classical and data-driven path planning methods against our proposed iA* framework through extensive experiments on public datasets and simulation environments. The experimental results show that iA* achieves an average of 6.3% better trade-off between path length and search area, while saving 11.5% runtime, 3.3% search areas, and shortening 5.6% path length than supervised learning methods for unseen maps, demonstrating its significantly better generalization abilities.
@@ II-A Classical Path Planning
The path planning problem, finding a collision-free path from the start node to the goal node with a given map, has been developed for many years. In the beginning, the classical works solve the path planning problem by graph-search approaches [west2001introduction], which iteratively search the nodes from the start node to the goal node, such as Breadth First Search (BFS). After that, some algorithms, such as Greedy BFS [doran1966experiments], A* [hart1968formal], and D* [stentz1994optimal], incorporate the heuristic information and provide a search direction to improve the planning efficiency. These heuristic search-based algorithms can find the optimal path when their suitable heuristic functions are admissible. However, these search-based methods face the problem of low search efficiency with the increasing size of maps due to their exponential time complexity growth.
Unlike search-based path planning methods, sampling-based methods sample nodes and connect nodes to find a solution in the given map. As the number of sampling points increases, the solution is closer to the optimal path. Typically, they stop their sampling when finding a near-optimal path acceptable in the given task to improve their search efficiency. The sampling-based methods can alleviate the computational growth by increasing the dimension of maps. By using a tree structure to store nodes and establish connections between the sampled points, some methods have shown outstanding performance, such as RRT [lavalle2001rapidly] and its derivatives [gammell2014informed, gammell2020batch, strub2022adaptively]. Besides, some sampling-based path planning methods [bohlin2000path, ravankar2020hpprm, liu2023prm] utilize the probabilistic road map to get solutions. However, these sampling-based methods can only get the near-optimal solution and may fail to find paths in complex environments with high obstacle density.
@@ II-B Learning-based Path Planning
Recently, with the development of deep learning technology, plenty of learning-based path planning methods have appeared. Some of them utilize networks to extract environmental information and provide guidance for improving search efficiency. SAIL [choudhury2018data] aims to find a suitable heuristic function that could observe and process the whole map instead of the local map during planning. Besides, [takahashi2019learning] and TransPath [kirilenko2023transpath] integrate the U-net-based model [ronneberger2015u] and Transformer-based model [vaswani2017attention] to learn their heuristic functions, respectively. Some methods [wang2020neural, yonetani2021nastar] utilize their well-trained network to encode the original map and generate a guidance map. Then, the guidance map allows them to reduce their search area and improve their search efficiency. MPNet[qureshi2020motion] uses two networks, Encoder net (Enet) and Planning net (Pnet), for the path planning task, where Enet handles the environment information, and Pnet can provide the next state of the robot under the current situation.
Some researchers use reinforcement learning (RL) [kaelbling1996reinforcement] to solve the path planning task. They aim to find an end-to-end solution for the path planning task without any policies. They formulate the path planning problem as a Markov decision process where the next state in the path depends on the current state and surroundings. [panov2018grid] demonstrates that it is possible to solve the path planning problem in grid maps using the RL-based method. [gao2019global] combines the path graph with Q-learning, efficiently finding the optimal path. However, these learning-based methods face problems of lacking interpretability and extensive data labeling.
@@ II-C Imperative Learning
Imperative learning (IL) [wang2025imperative] is an emerging self-supervised neuro-symbolic framework based on bilevel optimization, where the upper-level optimization is a data-driven model and the lower-level optimization is an interpretable reasoning engine. IL adopts a self-supervised manner. The results from lower-level optimization can provide an adaptive and precise gradient for the training of the data-driven model within upper-level optimization. For example, iPlanner [yang2023iplanner] utilizes a B-spline interpolation algorithm to guide the network training, generating safe and smooth trajectories with depth images. iSLAM [fu2023islam] adopts the IL framework to integrate the front-end and back-end of visual odometry. The pose estimation guides the training to generate the poses aligned with geometrical reality. iMatching [zhan2024imatching] utilizes the bundle adjustment module to train a neural model, generating the optimal translation matrix for pose estimation. iKap [li2025ikap] leverages a differentiable model predictive control to achieve a kinematics-aware vision-to-planning framework. These works demonstrate the efficiency of IL in robot navigation tasks. To the best of our knowledge, iA* is the first method using the IL to solve global path planning problems.
@@ III Methodology
Traditional search-based methods can guarantee finding the optimal paths, while deep neural networks have the ability to reduce the search area by extracting the map information and making predictions. To leverage the advantages of both learning-based and search-based path planning methods, the proposed imperative A* (iA*) adopts the imperative-learning (IL) strategy, including learning-based upper-level optimization and A* based lower-level optimization.
@@ III-A Framework
The framework of iA* is depicted in Fig. 2, which takes a planning instance as input, a three-layer tensor containing the information of obstacles, start, and goal positions, and outputs the near-optimal path (red) and search area (green). The iA* framework can be formulated as a bilevel optimization:
@@
Figure 2: The framework of iA* consists of three main components: (1) an upper-level optimization module featuring an instance encoder that processes the map, start, and goal, (2) a lower-level optimization module that integrates a differentiable A* search for node selection and expansion, and (3) a memory storage module that maintains intermediate representations, $H$, $G$, and $F$ to facilitate interaction between the upper and lower level optimizations.
The main difference between previous methods, like Neural A* [yonetani2021nastar], and iA* is that the bilevel optimization (1) incorporates a metric-based optimization to provide a more accurate training direction for the network $f$, leading to a more “generalizable” smaller search area. Intuitively, if the lower-level optimization (1b) is removed and $\mu^{*}$ represents constant optimal paths (labels), then bilevel optimization (1) degrades to a single-level optimization and iA* becomes Neural A*, and then the network gradients $\nabla_{\theta}\mathcal{U}$ are
@@
$\nabla_{\theta}U=\frac{\partial U}{\partial f_{\theta}}\frac{\partial f_{\theta}}{\partial\theta}+\boxed{\frac{\partial U}{\partial\mu^{*}}\frac{\partial\mu^{*}}{\partial L}({\frac{\partial L}{\partial f}\frac{\partial f}{\partial\theta}}+\frac{\partial L}{\partial\mu}+\frac{\partial L}{\partial M})},$ (3)
where the $\boxed{\text{second term}}$ illustrates an important difference. That means, the training of iA*, in the $t$-th iteration, relies on the optimal path $\mu_{t}$ for a search area given by $f(\theta_{t})$. Since the searching process defined by (1b) contributes to the gradient calculation, the network tends to predict a lower searching-cost area (lower-level optimization cost $L$) and thus more “likely” converge to better solutions. Nevertheless, this process requires a differentiable lower-level optimization, thus we next illustrate its details.
@@ III-B Lower-level Optimization
The lower-level optimization module is designed to achieve two objectives: (1) executing the searching process for path planning, and (2) providing differentiable computational gradients for back-propagation in upper-level network training. The classical A* algorithm inherently relies on a discrete graph structure, presenting challenges in effectively providing differentiable losses for upper-level optimization. To overcome this critical challenge, the lower-level optimization employs differentiable A* (dA*) [yonetani2021nastar] as its search engine, back-propagating the loss of the path planning process.
In dA*, all variables and parameters are represented in matrix formulations. Specifically, denote the open-list matrix $O$ and closed-list matrix $C$, which are binary and of the same size as the given map, to record the candidate nodes under evaluation and the nodes that have already been evaluated, respectively. The start node $n_{s}$, the goal node $n_{g}$, and the currently selected node $n^{\prime}$ are separately represented by one-hot indicator matrices $N_{s}$, $N_{g}$, and $N^{\prime}$, where only a single element is set to $1$, and all others are 0. Denote $s(n)$ as the accumulated path length from the start node and $h(n)$ as the estimated heuristic distance to the goal. The matrices $S$ and $H$ store the $s$ and $h$ values of each node. During the path planning process, dA* iteratively selects node $n^{\prime}$ with the minimum $s(n)+h(n)$ from the open list $\mathcal{O}$ and updates the $S$ matrix accordingly, until the optimal path is found.
In Neural A*, the neural network estimates the initial $s$ values for each node to accelerate the searching process. However, this manner results in error propagation: inaccuracies in $s(n)$ estimation during earlier iterations accumulate and affect subsequent search as the $s$ values are updated relying on the prior $s$. To eliminate this problem, we introduce a prediction factor $p$ in iA* to independently represent the influence of upper-level optimization in the Node Selection part of dA*. $P$ is the matrix storing the predicted $p$ values of each node. Then, we iteratively select the node with minimum $s(n)+h(n)+p(n)$ in the $\mathcal{O}$. It is worth noting that if the $p(n)$ equals $(\omega-1)h(n)$, our A* search algorithm equals weighted A* search [pohl1970heuristic], where the $\omega$ is a pre-set weight factor amplifying the influence of the heuristic function. In iA*, the node selection is formulated as:
@@
where $A\odot B$ is the Hadamard product [horn2012matrix], $\tau$ is an empirically defined parameter, the same as dA*, and $\mathcal{L}_{max}(A)$ is the function to obtain the argmaxargmax index matrix of $A$.
@@ III-C Upper-level Optimization
The upper-level optimization aims to increase the search efficiency and keep the sub-optimality. This objective also balances the search area and path length. We take advantage of a neural network in the upper-level optimization because of its great power in information extraction and prediction. The neural model is a widely-used U-Net [ronneberger2015u], where the decoder is used for feature extraction and the encoder for prediction. The use of U-Net ensures that the shape of the predicted $P$ is consistent with the shape of the input maps, avoiding changing the structure of the network and retraining in new environments with different shapes, and providing convenience in robot applications.
Figure 3: The calculation of the path length. $\mu$ is the Path Matrix generated from the lower-level optimization, AKA*K represents the convolution operation between matrix $A$ and kernel $K$, and $D$ is the distance matrix for further path length calculation.
@@
where $U(\mu)$ is the target function of upper-level optimization given $\mu$, $\mathcal{L}_{a}(\mu)$, $\mathcal{L}_{l}(\mu)$ are the search area loss and path length loss of path $\mu$, and $w_{a}$ and $w_{l}$ are the weights to adjust the ratio of the losses. We define $\mathcal{L}_{a}(\mu)$ as the number of extra visited nodes, which are visited by the algorithm but are not part of the final planned path, in one single path planning process. The calculation is shown as follows:
$\mathcal{L}_{a}(\mu)=\sum(C_{\mu}-\mu),$ (6)
@@
where the $w_{a}$ and $w_{l}$ are set to 1 empirically for keeping the optimal path while reducing the search area.
In iA*, the upper-level optimization updates parameters in the embedded network to predict the $P$, pursuing the balance between search area and path length while maintaining near-optimal path quality. Unlike the other data-driven methods [yonetani2021nastar, kirilenko2023transpath], directly calculating loss with constant optimal paths, our method dynamically calculates loss from the path planning results of lower-level optimization in each iteration. Thus, the network parameters are updated to find the optimal balance between search area and path length under temporary conditions, instead of invariably following the constant optimal paths. To validate this bilevel training process, we compare the effects of the conventional optimal-path-based training process with ours in the ablation study.